# Draft: August 22, 2025

### v1

@@ IV-B Case of Tropospheric $\text{CO}_{2}$
We now consider realistic initial condition and set-point to tackle the control of the concentration of tropospheric carbon. Specifically, we consider the current state of the atmosphere as initial conditions and the pre-industrial condition as the set-point state. In other words, the controller is expected to bring the concentration of $\text{CO}_{2}$ to the pre-industrial era.
Let $\bm{x}_{i0}$ be the initial condition of the $i$-th state. We consider a target urban area with volume $V_{1}=1.18\text{ km}^{3}$, which results from the product between the surface of the Italian city of Siena (approximately 118 $\text{km}^{2}$) and an high of 0.01 km. Given that the current concentration of carbon dioxide in the atmosphere is 431 ppm [2], the current mass of carbon inside a volume $V_{1}$ is 915.4 tonnes; thus, $x_{10}$ = 915.4 t. The surrounding area $c^{4}_{4,4}$ has a volume twice the one of $c^{1}_{1,1}$, and hence, $x_{40}$ = 1,830.8 t. To set the initial mass of fuel (diesel) in $c^{2}_{2,2}$, i.e., $x_{20}$, we consider $n_{\text{q}}=5,000$ vehicles with a tank of 50 liters each; this yields 42 kg of diesel per vehicle, and hence, $x_{20}$ = 210 t. The initial mass of gas inside pipelines and in one house is set as 50 kg, and hence, with the number of houses set as $n_{\text{h}}=10,000$, we have that $x_{30}$ = 500 t.
@@
Thus, the state matrix becomes
@@
$\theta_{1}$ = -0.7, and $\theta_{2}$ = -0.4. The eigenvalues of $\bm{A}$ are -0.3, 0, -0.7, and -0.4, while the pair ($\bm{A},\bm{B}$) is stabilizable. By choosing $\bm{R}_{1}$ = $\bm{I}_{4}$, the pair ($\bm{A},\bm{R}_{1}$) is observable. With $r_{2}=1$, the optimal gain becomes $\bm{K}$ = [-0.94, -0.67, -0.93, -0.65]. The response of the closed-loop system to the initial conditions specified previously and the control input are shown in Fig. 3. As visible, the concentration in the pre-industrial era $x_{1\text{e}}$ is reached in approximately 25 days. Note that, since $x_{2\text{e}}$ = $x_{3\text{e}}$ = 0, all the mass flow rates converge to zero with the exception of the flows between $c^{1}_{1,1}$ and $c^{4}_{4,4}$, i.e., the tropospheric regions are a closed system at equilibrium. This result explains why the paper title specifies that the network is “finite-time combustion-tolerant”: to reach the desired pre-industrial era condition inside the target area, i.e., to reach $x_{1\text{e}}$, this control scheme brings to zero the $\text{CO}_{2}$ emissions of the $n_{\text{q}}$ vehicles and the $n_{\text{h}}$ houses, i.e, the mass flow rates from $c^{2}_{2,2}$ and $c^{3}_{3,3}$ to $c^{1}_{1,1}$ and $c^{4}_{4,4}$. Specifically, this full-state control scheme tolerates the combustion taking place in 5,000 vehicles and in 10,000 house heating systems for approximately 6 days, which is the time it needs to bring $x_{2}(t)$ and $x_{3}(t)$ to zero.
@@
The dynamics of the key variables are reported in Fig. 4. Initially, circularity $\lambda(t)$ is positive because the extraction of carbon, i.e., $u(t)$, dominates over the other flows; at equilibrium, the extraction of finite natural resources $\phi_{1}(t)$ reaches zero, and so does the net-zero variable $\phi_{\text{nz}}(t)$. In particular, this controller meet the net-zero target (16) by nullifying both the input flows and the output flows of the tropospheric areas $c^{1}_{1,1}$ and $c^{4}_{4,4}$. Fig. 4 also shows the dynamics of the tropospheric temperature, for which we set the current atmospheric temperature reported by NASA [1] as the initial condition, i.e., $T(0)$ = 15.2 °C. The simulation is obtained for $C$ = 8 $\times$ 108 J K-1 m-2, $\alpha$ = 0.3, and $S$ = 1367.6 $\times$ 86400 J d-1 m-2 (values taken from [8]). The temperature has a dynamics much slower than that of the states since its first derivative starts to reduce after approximately 2000 days, that is, 5.5 years; after 16.4 years (6000 days), the temperature reaches the equilibrium value of 133 °C. This value is so high that it requires further investigation with climate experts of the model (17) (result similar to that obtained in [8, Fig. 7]). Finally, Fig. 4 shows the emissivity $\epsilon(t)$, which is inversely proportional to $\bm{x}_{1}(t)$.
Figure 4: Key variables with full-state feedback: (left) circularity, $\phi_{1}$, and $\phi_{\text{nz}}$; (middle) tropospheric temperature; (right) emissivity.
Remark: In terms of the implementation of (17) in combination with the closed-loop system, it is worth pointing out that the former is nonlinear, whereas the latter is linear. Thus, we achieved the dynamics of $T(t)$ by numerically integrating the resulting system of five ordinary differential equations with the Python function scipy.integrate.odeint(). Coherently with the states $x_{i}(t)|_{i=1,2,3,4}$, we translated the temperature to the condition in the pre-industrial era, and hence, $\tilde{T}(t)$ = $T(t)$ - $T_{\text{e}}$, with $T_{\text{e}}$ = 13.9 °C. However, $T_{\text{e}}$ may not be an equilibrium for (17), which is not required by the control law because the regulator acts on the states $x_{i}(t)|_{i=1,2,3,4}$, not directly on $T(t)$. The controller affects $T(t)$ indirectly through $\epsilon(t)$, which is a function of $x_{1}(t)$ in closed loop.
@@ V Numerical Studies: Output Feedback
This section covers numerical studies in the case of output feedback with the controller design detailed in Section III-D. Similarly to Section IV, first it tests the control with random rate constants $a_{k,s}$, and then, it considers and discusses the case of rate constants set to tackle the control of the tropospheric carbon and the carbon circularity in a target area $c^{1}_{1,1}$. For the output feedback, we leveraged the MATLAB implementations made available by Ilka and Murgovski [28].
V-A Random Rate Constants
We consider the same rate constants of Section IV-A, and hence, the matrix $\bm{A}$ corresponds to (38). Thus, the pair ($\bm{A},\bm{B}$) is stabilizable. The PBH test shows also that the pair ($\bm{A},\bm{C}$) is detectable. Similarly to [28], the weight matrices have been chosen as $\bm{R}_{1}=\bm{C}^{\top}\bm{C}$, $r_{2}=1$, and $\bm{N}=\bm{0}$. Thus, the condition (35) is satisfied. By running the Newton’s method and the modified Newton’s method of [28], we obtained $\bm{K}=-0.694$ with both algorithms. The two methods returned slightly different $\bm{P}$ matrices, but the difference between corresponding entries of the two returned $\bm{P}$ matrices were smaller than 10-7, and hence, both methods gave approximately the same $\bm{G}$ = [0, 0.696, 0.698, 0.618]. Furthermore, condition (36) was satisfied by both methods since the left-hand side of the equality contained entries smaller than 10-7. The implementation of the closed-loop system (22) with testing initial conditions $\bm{x}_{0}$ = [1, 2, 3, 4] and set-point state $\bm{x}_{\text{e}}$ = [1, 0, 0, 0.605] gave the results shown in Fig. 5. The controller stabilizes the translated state to the origin because its goal is to minimize the performance functional (24), while for the original state and control input we have that $\lim_{t\to\infty}\bm{x}(t)=\bm{x}_{e}$ and $\lim_{t\to\infty}u(t)=v_{\text{e}}$ (remember that $v_{\text{e}}=0$ from equation (30)). In addition, for the state in closed loop it holds that $\bm{x}(t)\geq\geq 0$, which is physically consistent.
Figure 5: Closed-loop state, output, and control input with output-feedback gain computed via the Newton’s method [28] in the case of random rate constants $a_{k,s}$.
@@ V-B Case of Tropospheric $\text{CO}_{2}$
This section addresses the control of tropospheric carbon as in Section IV-B, but designing an output-feedback control. Thus, the state matrix is that in equation (39) and the control design procedure follows that of Section V-A with matrices tuned as $\bm{R}_{1}=\bm{C}^{\top}\bm{C}$, $r_{2}=1$, $\bm{N}=\bm{0}$, and $\bm{R}_{1}=\bm{I}_{4}$. Hence, the pair ($\bm{A},\bm{B}$) is stabilizable, the pair ($\bm{A},\bm{C}$) is detectable, the condition (35) is satisfied, while the Newton’s and the modified Newton’s methods of [28] yield $\bm{K}$ = -0.837. The two methods returned slightly different $\bm{P}$ matrices, but the difference between corresponding entries of the two returned $\bm{P}$ matrices were smaller than 10-7, and hence, both methods gave approximately the same $\bm{G}$ = [0, 0.284, 0.357, 0.089]. Furthermore, condition (36) was satisfied by both methods since the left-hand side of the equality contained entries smaller than 10-7. The implementation of the closed-loop system with the initial and set-point conditions used in Section IV-B gives the results reported in Fig. 6. As visible, the concentration of $\text{CO}_{2}$ in the target area reaches the value of the pre-industrial era (i.e., 637.2 tonnes) in approximately 60 days, which is about 35 days slower than the system with full-state control.
Figure 6: Closed-loop state, output, and control input with output-feedback gain computed via the Newton’s method [28] for tropospheric $\text{CO}_{2}$ control.
The fact that the action of the output-feedback controller is slower than the full-state one is explained by comparing the control inputs: $u(0)\approx$ 260 t/d in the former and $u(0)\approx$ 1200 t/d in the latter, which is motivated by the fact that the full-state acts on all the four states, whereas the output-feedback only on one state (i.e., $x_{1}(t)$).
The key variables with output-feedback are shown in Fig. 7. In particular, $\lambda(0)$ is negative in contrast with the case of full-state feedback since, as discussed above, the control input $u(0)$ is more than 4 times smaller. Analogously to the case of full-state feedback, $\lambda(t)$, $\phi_{1}(t)$, and $\phi_{\text{nz}}(t)$ converge to zero, while the tropospheric temperature converges to 133 °C in approximately 16.4 years. In contrast, the emissivity $\epsilon(t)$ has a different response in the full-state setting compared to the output-feedback setting.
Figure 7: Key variables with output feedback: (left) circularity, $\phi_{1}$, and $\phi_{\text{nz}}$; (middle) tropospheric temperature; (right) emissivity.
@@ VI Conclusion
This paper proposes CarboNet, a network of thermodynamic compartments controlled via LQRs for tropospheric carbon regulation. The controllers effectively reduced the $\text{CO}_{2}$ concentration to the pre-industrial era level and also met the net-zero target, but they tolerated $\text{CO}_{2}$ emissions for only 6 days. The tropospheric temperature stabilizes at 133 °C in 16.4 years. The high of this value suggests to further investigate the temperature model in collaboration with climate experts, which is left as future work along with making the control system able to tolerate the combustion for more than 6 days.

### v2

@@ IV-B Case of Tropospheric $\text{CO}_{2}$
We now consider realistic system initial conditions and set-point to address the control of the concentration of tropospheric carbon. Specifically, we consider the current state of the atmosphere as our initial condition and the pre-industrial condition as the set-point state. In other words, the controller is expected to bring the concentration of $\text{CO}_{2}$ to the pre-industrial era.
Let $x_{i0}$ be the initial condition of the $i$-th state. We consider a target urban area with volume $V_{1}=1.18\text{ km}^{3}$, which results from the product between the surface of the Italian city of Siena (approximately 118 $\text{km}^{2}$) and a hight of 0.01 km. Given that the current concentration of carbon dioxide in the atmosphere is 431 ppm [NASAcarbon], the current mass of carbon inside the volume $V_{1}$ is 915.4 tonnes; thus, $x_{10}$ = 915.4 t. The surrounding area $c^{4}_{4,4}$ has a volume twice the one of $c^{1}_{1,1}$, and hence, $x_{40}$ = 1,830.8 t. To set the initial mass of fuel (diesel) in $c^{2}_{2,2}$, i.e., $x_{20}$, we consider $n_{\text{q}}=5,000$ vehicles with a gas tank of 50 liters each. This yields 42 kg of diesel per vehicle, and hence, $x_{20}$ = 210 t. The initial mass of gas inside pipelines and in one house is set as 50 kg, and hence, with the number of houses set as $n_{\text{h}}=10,000$, we have that $x_{30}$ = 500 t.
@@
Thus, the system matrix (9) becomes
@@
with $\theta_{1}$ = -0.7 and $\theta_{2}$ = -0.4. The eigenvalues of $\bm{A}$ are -0.3, 0, -0.7, and -0.4, and the pair ($\bm{A},\bm{B}$) is stabilizable. By choosing $\bm{R}_{1}$ = $\bm{I}_{4}$, the pair ($\bm{A},\bm{R}_{1}$) is observable. With $r_{2}=1$, the optimal gain is $\bm{K}$ = [-0.94, -0.67, -0.93, -0.65]. The response of the closed-loop system to the initial conditions specified previously and the control input are shown in Fig. 3. As can be seen, the concentration in the pre-industrial era $x_{1\text{e}}$ is reached in approximately 25 days. Note that, since $x_{2\text{e}}$ = $x_{3\text{e}}$ = 0, all the mass flow rates converge to zero with the exception of the flows between $c^{1}_{1,1}$ and $c^{4}_{4,4}$, i.e., the tropospheric regions are a closed system at equilibrium. This result explains why the paper title specifies that the network is “finite-time combustion-tolerant”: to reach the desired pre-industrial era condition inside the target area, i.e., to reach $x_{1\text{e}}$, this control scheme brings to zero the $\text{CO}_{2}$ emissions of the $n_{\text{q}}$ vehicles and the $n_{\text{h}}$ houses, i.e, the mass flow rates from $c^{2}_{2,2}$ and $c^{3}_{3,3}$ to $c^{1}_{1,1}$ and $c^{4}_{4,4}$. Specifically, this full-state control scheme tolerates the combustion taking place in 5,000 vehicles and in 10,000 house heating systems for approximately 6 days, which is the time needed to bring $x_{2}(t)$ and $x_{3}(t)$ to zero.
@@
The dynamics of the key variables are shown in Fig. 4. Initially, circularity $\lambda(t)$ is positive because the extraction of carbon, i.e., $u(t)$, dominates over the other flows; at equilibrium, the extraction of finite natural resources $\phi_{1}(t)$ reaches zero, and so does the net-zero variable $\phi_{\text{nz}}(t)$. In particular, this controller meets the net-zero target (15) by nullifying both the input flows and the output flows of the tropospheric areas $c^{1}_{1,1}$ and $c^{4}_{4,4}$. Fig. 4 also shows the dynamics of the tropospheric temperature, for which we set the current atmospheric temperature reported by NASA [NASAtemperature] as the initial condition, i.e., $T(0)$ = 15.2 °C. The simulation is obtained for $C$ = 8 $\times$ 108 J K-1 m-2 [elsherif2025climate], $\alpha$ = 0.3 [elsherif2025climate], and $S$ = 1361/4 $\times$ 86400 J d-1 m-2 [NASAirradiance]. The temperature has dynamics much slower than that of the states since its first derivative starts to reduce after approximately 5000 days, that is, 13.7 years; in the uncontrolled case (i.e., no carbon capture $u$ = 0), the temperature reaches the steady-state value of 35.2 °C in approximately 15000 days (41 years). In contrast, the steady-state value is 13.5 °C with full-state feedback control, which is 21.7 °C lower than the case of no carbon capture. Note that the steady-state temperature with closed loop (13.5 °C) is very close to the tropospheric temperature in the pre-industrial era (13.9 °C), which shows that the controller brings the temperature back to the pre-industrial era condition, as desired. Finally, Fig. 4 shows the emissivity $\epsilon(t)$, which is inversely proportional to $\bm{x}_{1}(t)$.
Figure 4: Key variables with full-state feedback: (left) circularity, $\phi_{1}$, and $\phi_{\text{nz}}$; (middle) tropospheric temperature with open and closed loops; (right) emissivity.
Remark: In terms of the implementation of (16) in combination with the closed-loop system, it is worth pointing out that the former is nonlinear, whereas the latter is linear. Thus, we achieved the dynamics of $T(t)$ by numerically integrating the resulting system of five ordinary differential equations with the Python function scipy.integrate.odeint(). Coherently with the states $x_{i}(t)|_{i=1,2,3,4}$, we translated the temperature to the condition in the pre-industrial era, and hence, $\tilde{T}(t)$ = $T(t)$ - $T_{\text{e}}$, with $T_{\text{e}}$ = 13.9 °C. However, $T_{\text{e}}$ may not be an equilibrium for (16), which is not required by the control law because the regulator acts on the states $x_{i}(t)|_{i=1,2,3,4}$, and not directly on $T(t)$. The controller affects $T(t)$ indirectly through $\epsilon(t)$, which is a function of $x_{1}(t)$.
@@ V Numerical Studies: Output Feedback
This section presents numerical studies for the output feedback case with the controller design detailed in Section III-D. As in Section IV, we first test the control with arbitrarily chosen rate constants $a_{k,s}$, and then, consider and discuss the case of rate constants set to tackle the control of the tropospheric carbon and the carbon circularity in a target area $c^{1}_{1,1}$. For the output feedback problem, we leveraged the MATLAB algorithm implementation made available by Ilka and Murgovski [ilka2022novel].
V-A Arbitrary Rate Constants
We consider the same rate constants given in Section IV-A, and hence, the system matrix $\bm{A}$ is given by (37). Thus, the pair ($\bm{A},\bm{B}$) is stabilizable. The PBH test shows also that the pair ($\bm{A},\bm{C}$) is detectable. As in [ilka2022novel], the weighting matrices have been chosen as $\bm{R}_{1}=\bm{C}^{\top}\bm{C}$ and $r_{2}=1$. Thus, condition (34) is satisfied. By running Newton’s method and the modified Newton’s method of [ilka2022novel], we obtained $\bm{K}=-0.694$ with both algorithms. The two methods returned slightly different $\bm{P}$ matrices, with the difference between corresponding entries of the two returned $\bm{P}$ matrices being smaller than 10-7, and hence, both methods gave approximately the same $\bm{G}$ = [0, 0.696, 0.698, 0.618]. Furthermore, condition (35) was satisfied by both methods since the left-hand side of the equality contained entries smaller than 10-7. The implementation of the closed-loop system (20) with $\hat{\bm{A}}=\bm{A}-\bm{BKC}$, testing initial conditions $\bm{x}_{0}$ = [1, 2, 3, 4], and set-point state $\bm{x}_{\text{e}}$ = [1, 0, 0, 0.605] gave the results shown in Fig. 5. The controller stabilizes the translated state to the origin because its goal is to minimize the performance functional (22), while for the original state and control input we have that $\lim_{t\to\infty}\bm{x}(t)=\bm{x}_{e}$ and $\lim_{t\to\infty}u(t)=v_{\text{e}}$ (recall that $v_{\text{e}}=0$ from equation (28)). In addition, the closed-loop system state $\bm{x}(t)\geq\geq 0,t\geq 0$, which is physically consistent.
Figure 5: Closed-loop state, output, and control input with output-feedback gain computed via the Newton’s method [ilka2022novel] in the case of arbitrary rate constants $a_{k,s}$.
@@ V-B Case of Tropospheric $\text{CO}_{2}$
This section addresses the control of tropospheric carbon as in Section IV-B with output-feedback control. Thus, the system matrix is given by (38) and the control design procedure follows that of Section V-A with $\bm{R}_{1}=\bm{C}^{\top}\bm{C}$ and $r_{2}=1$. Hence, the pair ($\bm{A},\bm{B}$) is stabilizable, the pair ($\bm{A},\bm{C}$) is detectable, and condition (34) is satisfied, while Newton’s and the modified Newton’s methods of [ilka2022novel] yield $\bm{K}$ = -0.837. The two methods returned slightly different $\bm{P}$ matrices, with the difference between corresponding entries of the two returned $\bm{P}$ matrices being smaller than 10-7, and hence, both methods gave approximately the same $\bm{G}$ = [0, 0.284, 0.357, 0.089]. Furthermore, condition (35) was satisfied by both methods since the left-hand side of the equality contained entries smaller than 10-7. The implementation of the closed-loop system with the initial and set-point conditions used in Section IV-B gives the results reported in Fig. 6. As can be seen, the concentration of $\text{CO}_{2}$ in the target area reaches the value of the pre-industrial era (i.e., 637.2 tonnes) in approximately 60 days, which is about 35 days slower than the system with full-state feedback control.
Figure 6: Closed-loop state, output, and control input with output-feedback gain computed via the Newton’s method [ilka2022novel] for tropospheric $\text{CO}_{2}$ control.
The fact that the action of the output-feedback controller is slower than the full-state one is explained by comparing the control inputs: $u(0)\approx$ 260 t/d in the former and $u(0)\approx$ 1200 t/d in the latter, which is motivated by the fact that the full-state feedback uses the knowledge of all four states, whereas output-feedback only one measured state (i.e., $x_{1}(t)$).
The key variables with output-feedback are shown in Fig. 7. In particular, $\lambda(0)$ is negative in contrast with the case of full-state feedback since, as discussed above, the control input $u(0)$ is more than 4 times smaller. As in the case of full-state feedback, $\lambda(t)$, $\phi_{1}(t)$, and $\phi_{\text{nz}}(t)$ converge to zero, while the tropospheric temperature behaves analogously to the case of full-state feedback discussed in Section IV-B. In contrast, the emissivity $\epsilon(t)$ has a different response in the full-state case compared to the output-feedback case.
Figure 7: Key variables with output feedback: (left) circularity, $\phi_{1}$, and $\phi_{\text{nz}}$; (middle) tropospheric temperature with open and closed loops; (right) emissivity.
@@ VI Conclusion
This paper proposes CarboNet, a network of thermodynamic compartments controlled via LQRs for tropospheric carbon regulation. The controllers effectively reduced the $\text{CO}_{2}$ concentration to the pre-industrial era level and also met the net-zero target, but tolerating $\text{CO}_{2}$ emissions for only 6 days. With the closed-loop control, the tropospheric temperature stabilizes approximately to the pre-industrial era condition, i.e., to 13.5 °C, which is 21.7 °C lower than the steady-state temperature achieved without carbon capture. Future research will focus on controller designs that are able to tolerate the combustion for more than 6 days as well as consider more accurate models of the dynamics of the tropospheric temperature to produce more accurate predictions.